OpenStudy (anonymous):

One pipe can fill a pond in 3 weeks, and second pipe can fill the pond in 5 weeks. However, evaporation and seepage can empty the pond in 10 weeks. If both pipes are used, how long will it take to fill the pond? I was hoping someone could help me also understand how to set up equations in this context also

OpenStudy (phi):

these problems are almost like rate * time = distance in this case rate * time = filled pond where rate is measured in ponds per week and time is measured in weeks

OpenStudy (anonymous):

So I'm thinking 3t+5t = 1?

OpenStudy (phi):

3t says: 3 times number of weeks rates are fractions (or ratios) example 10 miles per hour is 10 miles /1 hour here you have filled ponds per week

OpenStudy (phi):

in other words 3t does not make sense.

OpenStudy (anonymous):

Oh woops, I tried the equation (1/3+1/5)t=1 and it felt sort of correct but I'm not totally sure

OpenStudy (anonymous):

Or wait maybe (1/3+1/5)t-1/10=1

OpenStudy (phi):

close, 1/10 is also a rate. you should multiply by it by t

OpenStudy (phi):

notice if you fill a pond in 3 weeks, ⅓ ponds/week * 3 weeks = 1 pond that makes sense.

OpenStudy (anonymous):

Oh I see (1/3+1/5-1/10)t=1